A 0.1 kg hockey puck, initially moved at 5 m/s, slides a distance of 50 meters across a horizontal ice surface and comes to a stop due to some friction


1
Expert's answer
2021-02-04T14:38:49-0500

Find the coefficient of friction.


"F_f=ma\to\mu N=ma\to \mu mg=ma\to \mu=\frac{a}{g}"


"s=\frac{v^2-v_0^2}{2a}\to a=\frac{v^2-v_0^2}{2s}=\frac{5^2-0^2}{2\cdot 50}=0.25(m/s^2)"


"\mu=\frac{0.25}{9.8}=0.026" . Answer